In the question below the firm is purchasing a cash register system for $1,075,733.

As a consultant you were asked to evaluate a lease versus buy analysis on the purchase of a new front end cash register system for a clothing store called More Clothes.

The cost of the front end system was given in the first question.

The owners of the store will obtain a loan for 8% for the full amount of the system. The loan would be amortized over the 5-year life of the system with payments made at the end of each year. The system is classified as special purpose; hence it falls into the MACRS 3 year class with depreciation rates of (33%; 45%, 15% and 7%).

If the system is purchase a maintenance contract will be obtained for $30,000 payable at the start of each of the next 4 years. The firm will upgrade the system at the end of four years and sell the current system for $95,000.

As an alternative, ABC Leasing will write a 4 year lease on the system, including maintenance, for payments of $325,000 at the beginning of each year. The firms marginal federal plus state tax rate is 35%.

On an EXCEL spread sheet compute the net advantage to leasing.

Clearly note the following on the EXCEL spreadsheet:

a) Initial cost of owning

b) Year one operating cash flow of owning

c) Net sale price (if owned)

d) Cost of Owning

e) Cost of capital used to evaluate project

f) Operating cost in year 1 associated with leasing

g) Cost of Leasing

h) Net advantage (disadvantage) of leasing
Answer #1

Step-1 Estimation of cost of capital

Since there is no equity infusion involved and only Debt is used for the project, Cost of Capital= Cost of Debt

I) MACR 3 year class denotes payment of Principle at the end of 1.5, 2.5, 3.5 & 4.5 years. Hence the amortization schedule looks like this :

Period end 1 2 3 4 5
Amortization Schedule 33% 45% 15% 7%
Principle paid out                        -                354,992          484,080          161,360          75,301
Outstanding Loan         1,075,733              720,741          236,661            75,301                   -  

II) Interest amount calculation

Interest paid out at the end of each period= (Opening outstanding loan+closing outstanding loan)/2 x Interest rate

For Y-1 Interest paid= 1,075,733*8%

For Y-2 Interest paid= (1,075,733+720,741)/2 x 8% similarly for Y-2 to Y-5.

Interest payment schedule looks like below:

Period End 1 2 3 4 5
Interest Amount            (86,059)              (71,859)          (38,296)          (12,479)          (3,012)

Total Interest paid out = (211,704)

Total Cost of Debt = (1-corporate tax rate) *Total interest paid = (1-35%) x (211,704) = (137,607.77)

Cost of Debt in % = Total Cost of Debt/ Loan amount = 137,607.77/1,075,73 = 12.79%

This cost of Debt will be used in the later stage to calculate NPV and thus evaluation of the two options- Buying vs Leasing

Step-2 NPV of Buying vs Leasing

Buying Lease
Product Cost Maintenance Contract Total Cashflow
1         (1,075,733)            (30,000)          (1,105,733)          (325,000)
2            (30,000)                (30,000)          (325,000)
3            (30,000)                (30,000)          (325,000)
4                  95,000            (30,000)                  65,000          (325,000)
NPV             (984,657)          (970,896)

For NPV calculation 12.79% calculated in step-1 has been taken as the rate.

So it can be seen than with the buying-option, there is a cash outflow of $ 984,657 compared to $ 970,896 outflow in case of the lease-option.

Hence, Leasing is the right solution for the project with net advantage of $13,761